What must be the velocity of a spacecraft if 1 ℎ on its clock is to correspond to 1 ℎ + 1 𝑠 on a clock on


the earth?

1
Expert's answer
2022-02-13T12:12:41-0500

Explanation & Calculation


  • Just use the definition of the formula of time dialation.

"\qquad\qquad
\begin{aligned}
\small T&=\small \frac{T'}{\sqrt{1-\large\frac{v^2}{c^2}}}\\
\small 3601\,s&=\small \frac{3600\,s}{\sqrt{1-\large\frac{v^2}{c^2}}}\\
\small v&=\small 0.0236c
\end{aligned}"